So, the problem is:
2x + y = 1
-x + 2y = 7

&it says to solve the system, but I don't remember how! &the two problems are on top of each other

Answer by Alan3354(69443)

There's more than one way to do it.
2x + y = 1
-x + 2y = 7
--------------
2x + y = 1
-2x + 4y = 14 2nd eqn times 2
-------------- Add
5y = 15
y = 3
------
2x + 3 = 1
x = -1